40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6471     Entry time: Fri Mar 30 10:20:51 2012
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: locking last night 

I was trying to make the DRMI lock more robust.

Increasing the gains of the oplev on SRM helped a lot, but the lock is still not solid enough for measurements.

According to some line injection tests, the SRCL and MICH signals show up in AS55Q with almost the same amplitudes.

I tried to diagonalize the input matrix (particularly MICH-SRCL in AS55) based on the result of the line injection tests, but I ran out the time.

Work continues.

ELOG V3.1.3-